OpenStudy (anonymous):

A rectangle garden has a perimeter of 28 yards. the width of the garden is 6 yards less than its length. what is the area of the garden, in square yards?

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Solve the following for w and L :\[\{2(w+L)=28,w=L-6\} \] A=w*L=4*10=40 yd^2